OpenStudy (anonymous):

One of my favorite questions ever. Prove the Euler identity e^(2*pi*i) = 1

OpenStudy (anonymous):

Different approaches encouraged =D

OpenStudy (unklerhaukus):

\[e^{i\pi}=-1\] \[e^{2i\pi}=(-1)^2=1\]

OpenStudy (turingtest):

first prove e^(i*x)=cosx+isinx by using series expansions of sin and cos then plug in e^(pi*i) into that formula

OpenStudy (anonymous):

That's a nice approach, Turing. Now prove it to a high schooler. LOL.

OpenStudy (experimentx):

i^2 = -1 ... expand e using tylor's seris use ^^ this property to prove cosx+isinx <-- also in tylor's form and put the value of x, you will have proof.

OpenStudy (turingtest):

I learned that in high school! (granted it was towards the end)

OpenStudy (anonymous):

Haha really? Series expansions?

OpenStudy (turingtest):

Taylor expansions are in AP calc part II, so yeah...

OpenStudy (anonymous):

Well cool. =)

OpenStudy (turingtest):

you are not in the US I presume, where they have the "AP" program

OpenStudy (anonymous):

Haha no I am.

OpenStudy (experimentx):

yeah .. that will work. I am sure ... cos(x) = 1 - x2/2! + x4/4! - x6/6! + ... + xn.cos(n.pi/2)/n! <--- these minses comes form i sin(x) = x - x3/3! + x5/5! - x7/7! + ... + xn.sin(n.pi/2)/n! <-- also here but still there is i

OpenStudy (turingtest):

and we know the expansion of e^x add the above and...

OpenStudy (turingtest):

boom we see that e^(ix)=cosx+isinx

OpenStudy (anonymous):

I like this approach. It relies on more intuitive math. Begin with the function f(x) = e^(ix) f'(x) = ie^(ix) So that means that f'(x) = i*f(x) In the complex plane, a multiplication by i is the same as a 90 degree rotation.

OpenStudy (turingtest):

there is a really fast proof of e^(ix)=cosx+isinx in the wiki aritcle based on differential equations, but that is clearly above high-school level

OpenStudy (turingtest):

very nice :) true, it's more intuitive, but treating i as a constant is only defined for real numbers in differentiation isn't it?

OpenStudy (anonymous):



OpenStudy (anonymous):

Um... no. I is a constant.

OpenStudy (turingtest):

I've never taken complex analysis so I just wasn't sure you could use the same rules yes, i is a constant, I guess I wanted a more rigorous proof that it works the same though it's imaginary since that since rarely found in high school calc problems

OpenStudy (turingtest):

*since we rarely see i in high school calc

OpenStudy (anonymous):

So what that tells me is that, in the complex plane, f'(x) is always a 90 degree rotation of f(x), so the function's derivative is always orthogonal to the function, creating a circle! =D